OpenStudy (anonymous):

Please help! Which of the following is a solution of x2 - 7x = -5?

OpenStudy (anonymous):

\[you~~~~meant~~~~~~~~~~~~~~~~~~~~~~~~x^2 - 7x = -5\]

OpenStudy (anonymous):

which methods do you know? factoring, Quadratic Formula Completing the square

OpenStudy (anonymous):

I would suggest factoring.

OpenStudy (anonymous):

\[x^2 - 7x = -5\]\[x^2 - 7x + 5 = -5 + 5\]\[x^2 - 7x + 5 = 0\] No not factoring.

OpenStudy (anonymous):

Oh thank you :)

OpenStudy (anonymous):

Factoring isn't good here, try quadratic formula.

OpenStudy (anonymous):

ax2 + bx + c = 0 is this the quadratic formula?

OpenStudy (anonymous):